OpenStudy (anonymous):

Differentiate the following function y=x^11e^x dy/dx=?

OpenStudy (rogue):

Use the product rule\[\frac {d}{dx} \left[ uv \right] = u \frac {dv}{dx} + v \frac {du}{dx}\]

OpenStudy (anonymous):

11+x?

OpenStudy (rogue):

\[\frac {dy}{dx} = \frac {d}{dx} \left[ x^{11} e^x \right]\]\[\frac {dy}{dx} = x^{11}e^x + 11x^{10}e^x\]

OpenStudy (rogue):

Which you can factor out to\[\frac {dy}{dx} = x^{10} e^x (x + 11)\]

OpenStudy (rogue):

The derivative of e^x is just e^x.

OpenStudy (anonymous):

Thank you Rogue! That makes sense to me.

OpenStudy (rogue):

Your welcome :)